OpenStudy (anonymous):

Use the graph of y=f(1/x) to find lim (x->infinity) f(x) and lim (x->negative infinity) f(x) f(x)=xe^x

OpenStudy (anonymous):

Calculus is a wonderful thing.. T_T

OpenStudy (anonymous):

giv meh a sec ok :)

OpenStudy (anonymous):

Thanks a lot :D

OpenStudy (anonymous):

its a pleasure :)

OpenStudy (anonymous):

http://www.sagemath.org/calctut/inflimits.html mayb this will help

OpenStudy (zzr0ck3r):

we need the graph

OpenStudy (zzr0ck3r):

wow nm im blind...

OpenStudy (anonymous):

There is no graph

OpenStudy (anonymous):

ohhhh so u want a graph calculator

OpenStudy (anonymous):

That they give us

OpenStudy (anonymous):

No, what I mean is that I don't really understand what they are asking me to do

OpenStudy (zzr0ck3r):

it asks you to use the graph of \(f(\frac{1}{x})=(\frac{1}{x})e^{\frac{1}{x}}\)

OpenStudy (anonymous):

I dont understand how the graph of f(1/x) will help me find the other two limits

OpenStudy (anonymous):

how does tha graph look like

OpenStudy (anonymous):

do they giv u a pic of tha graph

OpenStudy (anonymous):

They don't show us. They want us to use a graphing calculator to find it. I can do that, I just don't understand what they want us to do with it

OpenStudy (anonymous):

I will take a picture of it

OpenStudy (anonymous):

hold on a min I think I have a graph calculater 2 do tha job :)

OpenStudy (anonymous):

let meh check first tho

OpenStudy (anonymous):

This is what the problem looks like: #49

jimthompson5910 (jim_thompson5910):

This may help. I used geogebra to graph.

OpenStudy (anonymous):

I appreciate it, but it doesn't really help much. There's only one function visible there. My main question is what they want me to do with the graph of f(1/x) to help me find the two limits

jimthompson5910 (jim_thompson5910):

You start with f(x) = x*e^x

jimthompson5910 (jim_thompson5910):

then you replace every copy of x with 1/x to get f(1/x) = (1/x)*e^(1/x)

jimthompson5910 (jim_thompson5910):

which is g(x) in that graph posted

OpenStudy (anonymous):

OH MY GOD, I AM SO STUPID xD

OpenStudy (anonymous):

I realized that just before I saw what you typed LOL

OpenStudy (anonymous):

Ok, so I have another question

OpenStudy (anonymous):

http://www.freemathhelp.com/factoring-calculator.php this page has calculus calculator on it :)

OpenStudy (anonymous):

hope it helps

OpenStudy (anonymous):

Why does the limit of f(x) as x approaches infinity equal the limit of f(1/x) as x approaches 0 from the right side? http://www.slader.com/textbook/9780132014083-calculus-graphical-numerical-algebraic-3rd-edition/76/exercises/49/

OpenStudy (anonymous):

giv meh a sec

OpenStudy (anonymous):

Thanks!!

jimthompson5910 (jim_thompson5910):

look at the graph of f(x) and g(x) what is the limit as x approaches infinity for f(x)?

OpenStudy (anonymous):

As x approaches infinity, the limit is infinity

OpenStudy (anonymous):

The basis of one-sided limits is that when a function jumps suddenly from one value to another, it often is not possible to describe the function's behavior with a single limit

OpenStudy (anonymous):

if that doesn't help here is a website http://www.sagemath.org/calctut/onesided.html

OpenStudy (anonymous):

Yeah but my question is why does the limit of the one function equal the limit of the other one What I mean is where it says based on the example above on this site http://www.slader.com/textbook/9780132014083-calculus-graphical-numerical-algebraic-3rd-edition/76/exercises/49/

OpenStudy (anonymous):

*based on the graph above

OpenStudy (anonymous):

When it says the two limits of f(x) equal limits of f(1/x)

OpenStudy (anonymous):

OH WAIT

jimthompson5910 (jim_thompson5910):

as x approaches infinity, x*e^x gets extremely large so f(x) approaches infinity as x approaches 0 from the right, (1/x)*e^(1/x) gets larger and larger, effectively approaching infinity So that's why \[\Large \lim_{x \to \infty}(f(x)) = \lim_{x \to 0^{+}}(g(x)) = \infty\]

OpenStudy (anonymous):

Yes! That is exactly what I was going to say!! Thank you so much!!

OpenStudy (anonymous):

I miss precalc and trig xD

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (anonymous):

I don't know this stuff sooo :/ srry bout takin u around tha block @TheWaffleMan149 :/

OpenStudy (anonymous):

Nah, don't worry! I really appreciate the help!

OpenStudy (anonymous):

its a pleasure :) ig :/ lol

OpenStudy (anonymous):

do I still get a waffle 4 hard wrk lol

OpenStudy (anonymous):

a minute agoDelete

OpenStudy (anonymous):

Whoops

OpenStudy (anonymous):

soooo can I have 1 pls *~*

OpenStudy (anonymous):

Actually, i have one more question

OpenStudy (anonymous):

One more question actually. Why does it only say limc(x) as x approaches negative infinity equal f(1/x) as x approaches 0 from the left? Doesn't f(1/x) also approach 0 when x approaches infinity?

OpenStudy (anonymous):

*limf(x)

OpenStudy (anonymous):

ummmm giv meh a sec

OpenStudy (anonymous):

Note that some functions, such as f(x) = x, do not tend to a nite value as x becomes very large, but instead grow to innity. Other functions, such as 1 x2 , become very large even for nite values of x. So we will dene innite limits to cover those cases:

OpenStudy (anonymous):

*define infinite limits

OpenStudy (anonymous):

does that help @TheWaffleMan149

OpenStudy (anonymous):

I appreciate the effort, but not really

OpenStudy (anonymous):

im sooooo srry idk nothing bout calculus :/ but I think ik some1 that does let meh get him 4 u

OpenStudy (anonymous):

Don't worry about it! I appreciate you trying to help me!!!

OpenStudy (anonymous):

r u sure u don't want meh 2 get another person 2 help u @TheWaffleMan149 :(

jimthompson5910 (jim_thompson5910):

So you're asking why \[\Large \lim_{x \to -\infty}(f(x)) = \lim_{x \to 0^{-}}(g(x))\] is true?

OpenStudy (anonymous):

Sorta. I know that is true, but I want to know why it doesn't also equal lim(g(x)) when x approaches infinity as well. On the graph it approaches the x axis in two places, so that is why I am confused that this only accounts for when it approaches 0 from the left

OpenStudy (anonymous):

If that makes sense at all

jimthompson5910 (jim_thompson5910):

As x approaches negative infinity, f(x) approaches 0 (you see this when you graph f(x) = x*e^x) As x approaches 0 from the left, g(x) approaches 0. This is shown in that graph I posted above. So that shows why \[\Large \lim_{x \to -\infty}(f(x)) = \lim_{x \to 0^{-}}(g(x))\] is true. They are both equal to 0.

jimthompson5910 (jim_thompson5910):

And before, you saw how as x approached infinity, g(x) approached zero so that means we have 3 different limits equal to 0

OpenStudy (anonymous):

But why isn't that included in the answer? They only address the one approaching 0 from the left

OpenStudy (anonymous):

It only says that limf(x) as x approaches infinity is equal to limf(1/x) as x approaches 0 from the left. It says nothing about both of those being equal to limc(1/x) as x approaches infinity

OpenStudy (anonymous):

Sorry I keep reiterating this, I just have a hard time being able to tell who understands what I am saying when it is online

OpenStudy (anonymous):

Is it because you only need to know that one of g(x)'s limits equal 0 to know that limf(x) as x approaches negative infinity equals 0?

jimthompson5910 (jim_thompson5910):

"It only says that limf(x) as x approaches infinity is equal to limf(1/x) as x approaches 0 from the left." that is false \[\Large \lim_{x \to \infty}(f(x)) =\infty\] while \[\Large \lim_{x \to 0^{-}} \left(f\left(\frac{1}{x}\right)\right) = 0\] So, \[\Large \lim_{x \to \infty}(f(x)) \neq \lim_{x \to 0^{-}} \left(f\left(\frac{1}{x}\right)\right)\]

jimthompson5910 (jim_thompson5910):

where does it say that? In the book? or solutions manual?

OpenStudy (anonymous):

This website that I am getting this from has the answers to my book.

OpenStudy (anonymous):

My book lists the answers, the website gives a little bit of info on how to find them

OpenStudy (anonymous):

It's not a true or false question

OpenStudy (anonymous):

The section of the website where it says "Based on the graph above" is what is confusing me. The second line of it

jimthompson5910 (jim_thompson5910):

So you're not sure about the notation on the second line?

OpenStudy (anonymous):

No, my confusion is that g(x) approaches 0 in two places, so why does the second line only acknowledge when it approaches it from the left side of 0. It also approaches it when x approaches infinity. Does that make any sense?

jimthompson5910 (jim_thompson5910):

I see. Well on the second line when they have 0^{-}, they mean "approach 0 from the left" So this is saying that as you approach x = 0 from the left, g(x) = f(1/x) is approaching 0. As to why they don't mention the limit \[\Large \lim_{x \to \infty} \left(f\left(\frac{1}{x}\right)\right) = 0\] I'm not sure

jimthompson5910 (jim_thompson5910):

Actually g(x) approaches 0 in three places: as x approaches positive infinity, negative infinity and 0 from the left.

OpenStudy (anonymous):

Yeah. Thats my big question. Why do they only mention the 0 front the left

OpenStudy (anonymous):

Do you think that is is because they only need one to prove that it does approach 0 and is therefore equal to lim f(x) when x approaches infinity?

jimthompson5910 (jim_thompson5910):

Ok I reread the solution that you linked and I see what they want

jimthompson5910 (jim_thompson5910):

They provide the graph of f(1/x) and you use this to find limits for f(x) (x approaching positive and negative infinity)

jimthompson5910 (jim_thompson5910):

as x approaches 0 from the right, f(1/x) approaches infinity When you approach 0 from the right on 1/x, this means 1/x is approaching positive infinity (x is on the right side of 0, so it's positive, 1/x is positive) So that's where they are getting that first line under where it says "based on the graph above"

jimthompson5910 (jim_thompson5910):

So that answers the first part of the question (which they show under the "result" portion)

OpenStudy (anonymous):

I understand that one. Its the second part that I don't understand. I don't understand why they ignored the other two places where the limit is 0

OpenStudy (anonymous):

I understand why they have what they do. My only confusion is why they left those out

jimthompson5910 (jim_thompson5910):

when x approaches 0 from the left, f(1/x) is approaching 0 x is now negative, so 1/x is growing to negative infinity therefore saying "x --> 0^{-}" is the same as "1/x ---> -infinity" so that's how they got that second line

jimthompson5910 (jim_thompson5910):

they are true facts, but they aren't needed to answer the original question "Use the graph of y=f(1/x) to find lim (x->infinity) f(x) and lim (x->negative infinity) f(x) f(x)=xe^x"

jimthompson5910 (jim_thompson5910):

we only care about what they are originally asking, which are the limits for f(x) x is approaching positive infinity x is approaching negative infinity

OpenStudy (anonymous):

So like I said earlier, it is just because they only need one of the 3 to prove that it has a limit at 0?

jimthompson5910 (jim_thompson5910):

I'm not sure how you can use the others, but yes, that is the key needed to prove the second line

OpenStudy (anonymous):

Ok. Thank you so much for the help!! I really, really appreciate it :D